# Hints of Prompt: Enhancing Visual Representation for Multimodal LLMs in Autonomous Driving

### v1

@@ 4.4 Ablation Study
To evaluate the contributions of each component within the HoP framework, we conduct comprehensive ablation studies. These include analyzing the effectiveness of each hint type independently (Fig. 5), assessing the impact of different hint sources (Tab. 4, Tab. 5, Tab. 6), and hyperparameters (Fig. 6). We also evaluate various fusion strategies for integrating the hints (Tab. 7) and, finally, analyze the efficiency gains of the proposed Efficient HoP model (Tab. 8).
@@
Sequential Cross-Attention 67.0 17.1 8.63
Parallel Cross-Attention 66.6 17.1 8.63
@@
Method #Params (M) Latency (ms)
LLaVA-v1.5 303.5 1535
HoP 832.0 2251
Efficient HoP 385.8 1548
Table 8: Efficiency of the Efficient HoP. #Params indicates the parameter count for all modules related to visual representation. Latency reports the average inference time measured on the LingoQA [34] validation set on an NVIDIA RTX 4090 GPU.
@@
Efficient HoP In the original HoP framework, we use DINOv2 [35] and Mask2Former [9] for Affinity and Semantic hints, which boost performance but increase computational costs. To reduce these, we create an Efficient HoP. As shown in Tab. 8, Efficient HoP introduces fewer additional parameters and latency compared to the HoP. Moreover, as shown in Sec. 3.3, Sec. 4, and Sec. 4, Efficient HoP surpasses the baseline across various metrics and achieves results close to those of the full HoP model.
@@ 5 Conclusion
In this paper, we introduce the Hints of Prompt (HoP) framework to address key limitations of current MLLMs in VQA tasks for autonomous driving. By incorporating Affinity, Semantic, and Question hints, HoP effectively captures instance-level structure, domain-specific elements, and query-specific context, leading to enhanced interpretability and reasoning capabilities. Extensive experiments on the LingoQA, DRAMA, and BDD-X datasets demonstrate that HoP consistently achieves state-of-the-art performance, validating the effectiveness of our proposed method. Additionally, we propose an Efficient HoP variant that reduces computational overhead while maintaining high performance. We believe that HoP’s structured integration of multi-level hints offers a promising direction for advancing VQA in complex, safety-critical scenarios.

### v2

@@ 4.4 Ablation Study
To evaluate the contributions of each component within the HoP framework, we conduct comprehensive ablation studies. These include data-efficient domain adaptation of HoP (Fig. 2), analyzing the effectiveness of each hint type independently (Fig. 6), assessing the impact of different hint sources (Tab. 4, Tab. 5, Tab. 6), and hyperparameters (Fig. 7). We also evaluate various fusion strategies for integrating the hints (Tab. 7) and, finally, analyze the efficiency gains of the proposed Efficient HoP (Tab. 8).
Figure 6: Impact of each hint type within the HoP. Abbreviations: AH (Affinity hint), SH (Semantic hint), QH (Question hint). “w/ cls” indicates the addition of class information to the Semantic hint, “ALL” denotes the combination of three hint types.
Data-efficient domain adaptation of HoP In Fig. 2, HoP trained with only 25% data achieves 64.0 Lingo-Judge score, surpassing the full-data performance of LLaVA-v1.5 (63.2), demonstrates remarkable data efficiency in driving-specific VQA task. More details are in the appendix.
@@
Parallel Cross-Attention 66.6 17.1 8.63
Sequential Cross-Attention 67.0 17.1 8.63
@@
Method LLM LJ $\uparrow$ #Params (M) Latency (ms)
VTS InternLM2-7B 64.2 304.0 740
LLaVA-v1.5 Vicuna-v1.5-7B 63.2 303.5 631
HoP Vicuna-v1.5-7B 67.8 832.0 956
Efficient HoP Vicuna-v1.5-7B 66.8 385.8 661
Qwen-v2.5-3B 64.6 385.8 504
Qwen-v2-0.5B 57.6 385.8 302
with AWQ [25] Qwen-v2.5-3B 64.2 385.8 281
Table 8: Performance and efficiency of Efficient HoP. #Params indicate the parameter count for all modules related to visual representation. Latency reports the average inference time measured on the LingoQA [35] validation set on an NVIDIA RTX 4090 GPU. AWQ [25] represents the quantized model.
@@
Efficient HoP In the original HoP framework, we use DINOv2 [36] and Mask2Former [9] for Affinity and Semantic hints, which boost performance but increase computational costs. To reduce these, we create an Efficient HoP. As shown in Tab. 8, Efficient HoP reduces parameters and latency compared with HoP while achieving comparable performance (Tab. 1, Tab. 2, and Tab. 3). The quantized Efficient HoP with the 3B model maintains this advantage with deployment-ready 281ms latency.
@@ 5 Conclusion
In this paper, we introduce the Hints of Prompt (HoP) framework to enhance the visual representation of MLLM in VQA tasks for autonomous driving scenarios. By incorporating Affinity, Semantic, and Question hints, HoP effectively captures instance-level structure, domain-specific elements, and query-specific context, enhancing interpretability and reasoning capabilities, and enabling rapid domain adaptation with limited data. Extensive experiments on the LingoQA, DRAMA, and BDD-X datasets demonstrate that HoP consistently achieves state-of-the-art performance. Moreover, we propose an Efficient HoP variant that reduces computational overhead while maintaining high performance. We believe that HoP’s structured integration of multi-level hints offers a promising direction for advancing VQA in complex, safety-critical scenarios.
Acknowledgement
This work was supported by the National Natural Science Foundation of China under Grant No. 62271466 and the National Key Research and Development Program of China Grant No. 2023YFF0716500.